# Minimizing resource overhead in fusion-based quantum computation using hybrid spin-photon devices

### v1

@@ Abstract
We present three schemes for constructing a (2,2)-Shor-encoded 6-ring photonic resource state for fusion-based quantum computing, each relying on a different type of photon source. We benchmark these architectures by analyzing their ability to achieve the best-known loss tolerance threshold for fusion-based quantum computation using the target resource state. More precisely, we estimate their minimum hardware requirements for fault-tolerant quantum computation in terms of the number of photon sources to achieve on-demand generation of resource states with a desired generation period. Notably, we find that a group of 12 deterministic single-photon sources containing a single matter qubit degree of freedom can produce the target resource state near-deterministically by exploiting entangling gates that are repeated until success. The approach is fully modular, eliminates the need for lossy large-scale multiplexing, and reduces the overhead for resource-state generation by several orders of magnitude compared to architectures using heralded single-photon sources and probabilistic linear-optical entangling gates. Our work shows that the use of deterministic single-photon sources embedding a qubit substantially shortens the path toward fault-tolerant photonic quantum computation.
@@ I Introduction
Figure 1: Stages of fusion-based quantum computation. Single photons can be combined using linear optics and measurements to produce entangled seed states, such as a linear-cluster state or a Greenberger–Horne–Zeilinger (GHZ) state. The black lines connecting photons indicate that a control-Z operation (CZ) was applied to the two photonic qubits each initially prepared in the |+=(|$\ket{+}=(\ket{0}+\ket{1})/\sqrt{2}$ state. States which can be described using this representation are called graph states. Multiple seed states can be combined to construct a larger graph state, the target resource state, that is sent to a fusion network to implement fusion-based quantum computation. This work focuses on the stages up to and including the target state production.
In this work, we study the steps required to construct resource states for FBQC (see Figure 1). We present three different architectures for constructing the 24-photon Shor-encoded (2,2) 6-ring resource state [6] that can achieve a loss tolerance of 7.5% when used in conjunction with biased measurements [16]. The first architecture is all-photonic, requiring only a source of single photons, either heralded or deterministic. The second uses a deterministic source of caterpillar graph states [17]. The third uses a group of such sources that can be entangled using repeat-until-success (RUS) control-Z gates [18].
@@ II.1.1 Heralded single-photon source
A heralded single-photon source (HSPS) is a single-photon source (SPS) that succeeds in the generation of a single photon with probability $p_{s}<1$ when triggered, and provides a heralding signal when it does succeed. This can be achieved, for example, by exploiting four-wave mixing in a material with a third-order nonlinearity [19]. Upon excitation by a strong pump laser pulse, two laser photons in the pulse may spontaneously annihilate leading to the creation of a pair of photons with a probability $p_{\textrm{pair}}$, called the idler and the signal, with different frequencies that sum to the laser frequency (see Figure 2a). The idler and signal can then be split into two different spatial modes so that the detection of the idler heralds the presence of its twin photon in the other spatial mode. Taking into account the heralding efficiency $\eta_{H}$, which is the probability that a photon is present in the signal mode when a photon was detected in the idler mode, and the idler detection efficiency $\eta_{\mathrm{det}}$, the probabilities $p_{s}$ and $p_{\textrm{pair}}$ are related by $p_{s}=\eta_{H}\eta_{\mathrm{det}}p_{\textrm{pair}}$.
Increasing the pump power improves the success probability $p_{s}$, but, when the pump is too intense, it becomes very likely that unwanted multiple pairs are produced instead. This leads to a multi-photon error probability that is proportional to $p_{s}$. Thus, $p_{s}$ is commonly limited to below $1\%$. One technique proposed to overcome this issue relies on photon-number resolving detectors to measure the idler and discard multi-photon events [20, 21]. In the limit of perfect heralding efficiency and photon-number resolution fidelity, this technique can boost the success probability up to at most $p_{s}=1/4$ [22], provided that the heralding efficiency is sufficiently high (eg. $\eta_{H}>0.99$) to reach the desired noise suppression (see Supplementary).
@@ II.1.2 Deterministic single-photon source
A deterministic SPS (DSPS) does not suffer from a fundamental trade-off between emission probability and multi-photon errors. It produces a nearly perfect single photon every time it is triggered. It has a success probability of $p_{s}=1$, and any missing photon is attributed to the difficulty of collecting the photon. Later on, we include the photon collection efficiency in the transmission efficiency (see III.1.3).
@@ II.1.3 Deterministic source of caterpillar graph states
Caterpillar graph states are useful building blocks for constructing resource states. Since a caterpillar source can also produce GHZ states on demand, fusion gates can be boosted to a success probability of $1-1/2^{b}$, where the positive integer $b\geq 2$ is the level of boosting, while consuming $22^{b}-$ 2 additional auxiliary photons per gate [40]. Alternatively, $2b$ additional dangling photons can be added to the building block during generation to allow repeated fusion attempts without using auxiliary states, again bringing the success probability up to $1-1/2^{b}$ [41]. This latter approach requires adaptivity, which can increase the chance that remaining dangling photons in the resource state are lost. In addition, to ensure a high success probability, many dangling photons must be produced even if the first fusion attempt ends up being successful. Since the gate will succeed after two attempts on average, many unused photons must be measured to decouple them from the resource state.
@@ II.1.4 Repeat-until-success module
To address the challenges with repeated fusions described above, an alternative is to entangle two or more caterpillar sources containing a qubit before emitting the fused graph (see Figure 2d for an example using two caterpillar sources). To entangle two sources, each source emits a photon that is sent to a small linear-optical module, where the two photons interact and get measured to implement a fusion operation. When one of the correct detection patterns is observed, this results in a controlled-Z operation applied on the spin qubits in each source. The operation is thus probabilistic, but it can be repeated until it succeeds [18]. In the absence of photon loss, the success probability $p_{s}$ can then arbitrarily approach 1 at the cost of taking more time to attempt gates.
This so-called RUS gate, and related RUS measurements, have recently been of great interest for designing hybrid spin-optical quantum computing (SPOQC) architectures tailored to quantum emitters containing a spin qubit [8, 9, 10], showing impressive thresholds. Here we demonstrate that they can also substantially improve a standard FBQC architecture, as the entanglement between the sources carries over to the emitted photons, thus allowing the generation of graph states with a more complex topology than caterpillar graphs, directly at source. Unlike repeated fusion gates, after entanglement is established, no additional dangling photons need to be produced. Therefore this approach ensures that a minimum number of photons are consumed to establish a graph edge. Moreover, the sources serve as a lossless quantum memory to store entanglement until other states are prepared.
@@ II.2.1 All-photonic scheme
In Figure 3a, we present an all-photonic scheme using six multiplexing stages. We designed this particular scheme to reduce the number of states needed to attempt a new stage while also maintaining a reasonable success probability for each stage. This provides a good compromise between boosting the success probability while minimizing the number of states being routed between stages. The construction begins with single photons, produced either probabilistically or deterministically. Four photons are combined to produce an intermediate 3-photon state, called a 2-primate, which succeeds with a probability of $\sim 28.4\%$ [44]. Two of these 2-primates can then be combined into a 4-photon GHZ states with a success probability of $\sim 9.1\%$ by measuring two of the photons (see Supplementary). From there, twelve 4-GHZ states are fused together in three consecutive stages to produce the target resource state.
Normally each unboosted fusion gate is expected to succeed with a probability of $50\%$. Here, to improve the probability of successfully generating the state, we assume that each gate is boosted to reach a success probability of $75\%$ using four additional single photons (not shown in Fig. 3a) [44]. Indeed, in the low-loss regime considered in our study, the cost of producing additional photons is lower than the cost of multiplexing more fusion gates. However, boosting beyond 75% requires additional GHZ states, which are themselves costly to produce from single photons and do not justify the mild increase to success probability compared to increasing the amount of multiplexing.
@@ II.2.2 Caterpillar scheme
The caterpillar scheme first applies two type-II fusion gates and three Y measurements to a 17-photon caterpillar graph to construct one-third of the resource state. The full 24-photon ring is then constructed by fusing three copies of those building blocks using three fusion gates (see Figure 3b for the fusion operations and exact caterpillar state required).
Although only a single caterpillar source is needed to produce the initial 17-photon state, the following stages require fusion gates. Thus, in this architecture, while substantially reduced, multiplexing is still required to overcome the probabilistic nature of fusion operations. As before, to optimize the success probability, all fusion gates are boosted to a success probability of $75\%$, but this time using a Bell pair (not shown in Fig. 3b) [40]. Note that a caterpillar source can also be used to efficiently boost fusion gates beyond 75%, however we find that the relative improvement compared to once-boosted gates is minor (see Supplementary) and may not justify the additional losses due to the increased complexity. Hence, we focus on the once-boosted fusion gate in our analysis.
@@ II.2.3 RUS scheme
The last approach that we introduce involves using RUS gates to entangle a small group of caterpillar sources. There are various ways to construct the resource state using interacting emitters [45, 46, 47, 48]. Here we propose a scheme aimed primarily at minimizing the optical depth for the generated resource state. This is achieved by first creating all the necessary entanglement links between the sources and only emitting the photonic state once this first step is successful. The entanglement between the sources is directly transferred to the photons generated in this second stage. In that case, photon loss occurs only due to the source itself and a single active switch is needed to release the photon from the RUS module (Figure 2d).
@@
Naively generating all 32 edges of the resource state requires the detection of 128 photons on average. Losing just one photon can destroy many edges and substantially increase the generation time. This problem is alleviated by recognizing that the 12 dangling photons of the resource state can be produced deterministically, provided that the highly-connected central ring of 12 qubits is first prepared across 12 sources. Furthermore, using the Graph State Compass (GSC) python package [49], we find that this smaller graph has a minimum-edge representation with just 14 edges, which can be transformed into the desired 12-qubit central ring using local complementation (LC) [^1] LC is a graph operation that corresponds to complementing the neighborhood of a vertex in a graph, and when two graphs can be obtained from one another by LC, the corresponding graph states can be obtained from one another by local Clifford operations [50]., see Figure 3c. The minimum-edge representation is hereafter called the LC-graph. This approach works only because the dangling photons can be produced afterwards, since the full 24-photon resource state is already in a minimum-edge representation.
@@
The LC-graph has further useful properties. Most qubits have a low degree of connectivity and so photon loss is less destructive. In addition, it can be constructed by connecting together identical subgraphs in three stages. Thus, if a photon is lost when generating one subgraph, it will not destroy the others.
@@
Once the LC-graph is complete, it is transformed into the central 12-qubit ring of the resource state using LC. Each source then emits a photon, mapping the prepared graph onto the emitted photons. Since the sources remain entangled to the photonic graph, they can emit a second photon to produce the 12 dangling photons of the final resource state. Finally each spin is measured to decouple it from the graph and reset for the next cycle.
@@ III.1 Figures of Merit
To compare architectures, we define two figures of merit quantifying the performance of an RSG: the resource efficiency $\eta_{R}$ and the transmission efficiency $\eta_{T}$. These metrics are designed to describe RSG architectures that operate on a periodic clock cycle with rate $r$, where a single resource state is produced by the RSG at the end of every cycle with a success probability $p_{\text{rsg}}$. This requirement is important to ensure synchronization between different RSGs and synchronization with the clock cycle of the fusion network required to perform FBQC.
@@ III.1.1 Resource efficiency
To benchmark RSG overhead, a platform-agnostic metric is needed that captures trade-offs across RSG architectures. This metric could encompass factors like the energetic cost of operation, the fabrication cost, or the spatial footprint of the device. Designing an inclusive and comprehensive definition of resource efficiency is a challenge that should be addressed in a future work. Here, we consider a simple definition that enables the comparison of RSG architectures exploiting spatial and temporal multiplexing schemes.
To quantify the resource efficiency in the spatial domain, we use the ratio of the number of photons in the final resource state $N$ to the number of sources $N_{0}$ used to produce at least one such resource state with a probability greater than a predefined value. In the temporal domain, we consider the ratio of the RSG clock rate $r$ to the rate $r_{0}$ of the laser pump used trigger the sources (the internal clock of the RSG). Combining these, we define a spatio-temporal resource efficiency of:
@@ III.1.2 Upper bound on $\eta_{R}$
A spatial MUX does not decrease RSG rate ($r=r_{0}$) and so $\eta_{R}=N/N_{0}$ depends only on the number of sources $N_{0}$. If the architecture must consume an average of $N_{\text{avg}}$ photons to produce one resource state, then a necessary condition to arbitrarily improve the RSG success probability $p_{\text{rsg}}$ toward $1$ is that the $N_{0}$ sources produce a minimum of $N_{\text{avg}}$ photons on average. Thus we have $p_{s}N_{0}\geq N_{\text{avg}}$ which implies $\eta_{R}\leq p_{s}N/N_{\text{avg}}=\tilde{\eta}_{R}$. This upper bound can be generalized to $r\neq r_{0}$ by noting that an optimal temporal MUX allows one to reduce $N_{0}$ at the cost of reducing $r$ by the same proportion. Hence:
@@
Knowing the upper bound, it is convenient to define the discrepancy between $\eta_{R}$ and $\tilde{\eta}_{R}$, which is directly attributed to the ability for a MUX to use the available photons. This leads to a natural definition of MUX efficiency $\eta_{\text{mux}}=N_{\text{avg}}r/N_{0}r_{0}$, so that
@@
This final expression is convenient as it clearly separates three independent contributions to the resource efficiency: the success probability $p_{s}$ that depends only on the source, the MUX efficiency $\eta_{\text{mux}}$ that depends only on the multiplexing network, and the efficiency of photon consumption $N/N_{\text{avg}}$ that depends only on the scheme.
The above derivation of $\eta_{\text{mux}}$ assumes that sources operate continuously for the entire RSG cycle. This is not the case for the RUS module because sources may idle after generating an edge. For example, in Figure 3c, the last operation before LC will be a RUS gate to establish the edge connecting just two of the 12 sources. When this RUS gate is being applied, the 10 other sources must wait without emitting photons, which increases the number of internal cycles per source. Thus $N_{\text{avg}}$ will underestimate the cost in time and overestimate the upper bound on resource efficiency. In this case, the upper bound can instead be computed knowing the average number of internal clock cycles $\tau_{\text{avg}}$ used to produce a resource state:

### v2

@@ Abstract
We present three schemes for constructing a (2,2)-Shor-encoded 6-ring photonic resource state for fusion-based quantum computing, each relying on a different type of photon source. We benchmark these architectures by analyzing their ability to achieve the loss tolerance threshold for fusion-based quantum computation using the target resource state. More precisely, we estimate their minimum hardware requirements for fault-tolerant quantum computation in terms of the number of photon sources to achieve on-demand generation of resource states with a desired generation period. Notably, we find that a group of 12 deterministic single-photon sources containing a single matter qubit degree of freedom can produce the target resource state near-deterministically by exploiting entangling gates that are repeated until success. The approach is fully modular, eliminates the need for lossy large-scale multiplexing, and reduces the overhead for resource-state generation by several orders of magnitude compared to architectures using heralded single-photon sources and probabilistic linear-optical entangling gates. Our work shows that the use of deterministic single-photon sources embedding a qubit substantially shortens the path toward fault-tolerant photonic quantum computation.
@@ I Introduction
[…]
In this work, we study the steps required to construct resource states for FBQC (see Figure 1). We present three different architectures for constructing the 24-photon Shor-encoded (2,2) 6-ring resource state [6] that can achieve a loss tolerance of 7.5% when used in conjunction with biased measurements [16]. The first architecture is all-photonic, requiring only a source of single photons, either heralded or deterministic. The second uses a deterministic source of caterpillar graph states [17]. The third uses a group of such sources that can be entangled using repeat-until-success (RUS) control-Z gates [18]. This comparative analysis, which focuses on resource overhead and loss tolerance at the level of state generation, constitutes a critical first step to assess the main advantages provided by emitters for FBQC.
@@ II.1.1 Heralded single-photon source
A heralded single-photon source (HSPS) is a single-photon source (SPS) that succeeds in the generation of a single photon with probability $p_{s}<1$ when triggered, and provides a heralding signal when it does succeed. This can be achieved, for example, by exploiting four-wave mixing in a material with a third-order nonlinearity [19]. Upon excitation by a strong pump laser pulse, two laser photons in the pulse may spontaneously annihilate leading to the creation of a pair of photons with a probability $p_{\textrm{pair}}$, called the idler and the signal, with different frequencies that sum to the laser frequency (see Figure 2a). The idler and signal can then be split into two different spatial modes so that the detection of the idler heralds the presence of its twin photon in the other spatial mode. Taking into account the heralding efficiency $\eta_{H}$, which is the probability of detecting the idler given that a pair was produced, the probabilities $p_{s}$ and $p_{\textrm{pair}}$ are related by $p_{s}=\eta_{H}p_{\textrm{pair}}$ (for ease of reading, all the parameters we introduce and their associated symbols are summarized in Appendix A).
Increasing the pump power improves the success probability $p_{s}$, but, when the pump is too intense, it becomes very likely that unwanted multiple pairs are produced instead. This leads to a multi-photon error probability that is proportional to $p_{s}$. Thus, $p_{s}$ is commonly limited to below $1\%$. One technique proposed to overcome this issue relies on photon-number resolving detectors to measure the idler and discard multi-photon events [20, 21]. In the limit of perfect heralding efficiency and photon-number resolution fidelity, this technique can boost the success probability up to at most $p_{s}=1/4$ [22], provided that the heralding efficiency is sufficiently high (eg. $\eta_{H}>0.99$) to reach the desired noise suppression (see Appendix C).
@@ II.1.2 Deterministic single-photon source
A deterministic SPS (DSPS) does not suffer from a fundamental trade-off between emission probability and multi-photon errors. It produces a nearly perfect single photon when triggered and does not require a heralding signal. As such, it has a success probability of $p_{s}=1$, and any inefficiency is treated as photon loss and absorbed into the probability of collecting the photon from the source. Later on, we also include the photon collection efficiency as part of the total transmission efficiency (see III.1.3).
@@ II.1.3 Deterministic source of caterpillar graph states
Caterpillar graph states are useful building blocks for constructing resource states. Since a caterpillar source can also produce GHZ states on demand, fusion gates can be boosted to a success probability of $1-1/2^{b}$, where the positive integer $b\geq 2$ is the level of boosting, while consuming $2^{b}-2$ additional auxiliary photons per gate [40]. Alternatively, $2b$ additional dangling photons can be added to the building block during generation to allow repeated fusion attempts without using auxiliary states, again bringing the success probability up to $1-1/2^{b}$ [41]. This latter approach requires adaptivity, which can increase the chance that remaining dangling photons in the resource state are lost. In addition, to ensure a high success probability, many dangling photons must be produced even if the first fusion attempt ends up being successful. Since the gate will succeed after two attempts on average, many unused photons must be measured to decouple them from the resource state (See Appendix B for details on the fusion gates).
@@ II.1.4 Repeat-until-success module
To address the challenges of repeated fusions described above, an alternative is to entangle two or more caterpillar sources containing a qubit before emitting the fused graph (see Figure 2d for an example using two caterpillar sources). To entangle two sources, each source emits a photon that is sent to a small linear-optical module. When one of the correct detection patterns is observed, this results in a controlled-Z operation applied on the spin qubits in each source. The operation is thus probabilistic, but it can be repeated until it succeeds [18]. In the absence of photon loss, the probability of successfully producing an entangled state can then arbitrarily approach 1 at the cost of taking more time to attempt gates.
This so-called RUS gate, and related RUS measurements, have recently been of great interest for designing hybrid spin-photon platforms for FTQC, such as the spin-optical quantum computing (SPOQC) architecture [8, 9] or the quantum-emitter-based FBQC scheme [10], showing impressive thresholds. Here, we demonstrate that it can also substantially improve a standard FBQC architecture, as the entanglement between the sources carries over to the emitted photons, thus allowing the generation of graph states with a more complex topology than caterpillar graphs, directly at source. Unlike repeated fusion gates, after entanglement is established, no additional dangling photons need to be produced. Therefore this approach ensures that a minimum number of photons are consumed to establish a graph edge. Moreover, the sources serve as a lossless quantum memory to store entanglement until other states are prepared.
In our illustrations and discussions we will use charged quantum dots as a typical example to implement caterpillar sources and RUS modules. However, our study is also relevant for other emitters containing at least one qubit degree of freedom, such as T centers [42], SnV centers [43], and atoms [35], to name a few.
@@ II.2.1 All-photonic scheme
In Figure 3a, we present an all-photonic scheme using six multiplexing stages. We designed this particular scheme to reduce the number of states needed to attempt a new stage while also maintaining a reasonable success probability for each stage. This provides a good compromise between boosting the success probability while minimizing the number of states being routed between stages. The construction begins with single photons, produced either probabilistically or deterministically. Four photons are combined to produce an intermediate 3-photon state, called a 2-primate [46]. Two of these 2-primates can then be combined into a 4-photon GHZ states by measuring two of the photons. From there, twelve 4-GHZ states are fused together in three consecutive stages of multiplexing to produce the target resource state. Each resource produced during each of these final three stages requires two fusion gates to succeed simultaneously. We also consider the use of additional auxiliary photons to increase the success probability of these fusion gates (see Section III.2).
@@ II.2.2 Caterpillar scheme
The caterpillar scheme first applies two type-II fusion gates to a 14-photon caterpillar graph to construct one-third of the resource state. The full 24-photon ring is then constructed by fusing three copies of those building blocks using three fusion gates (see Figure 3b for the fusion operations and exact caterpillar state required).
Although only a single caterpillar source is needed to produce the initial 14-photon state, the following stages require fusion gates. Thus, in this architecture, while substantially reduced, multiplexing is still required to overcome the probabilistic nature of fusion operations. As before, to optimize the success probability, we consider boosted fusion gates, but this time using a Bell pair (not shown in Fig. 3b), i.e. using Grice’s scheme [40], because a caterpillar source can generate Bell pairs deterministically with no overhead.
@@ II.2.3 RUS scheme
The last approach that we introduce involves using RUS gates to entangle a small group of caterpillar sources. There are various ways to construct the resource state using interacting emitters [47, 48, 49, 50, 51]. Here we propose a scheme aimed primarily at minimizing the optical depth for the generated resource state. This is achieved by first creating all the necessary entanglement links between the sources and only emitting the photonic state once this first step is successful. In this way, the probability of losing a photon in the resource state is only impacted by the efficiency of the source and the transmission of a single active switch (Figure 2d). This active switch is necessary to switch between the RUS entangling stage and the emission stage. Losses due to the entangling optics for the RUS gate only impact generation efficiency rather than the final resource state.
We apply CZ gates between the emitters using the RUS module described in Section II.1.4: each emitter sends a photon to a linear-optical interferometer, which implements a probabilistic rotated type-II fusion operation; if the fusion fails, we repeat the gate attempt, until we observe one of the correct detection patterns. Once the entanglement is built between the sources, it is directly transferred to the photons generated in the second stage (see Appendix B for details).
@@
Naively generating all 32 edges of the resource state requires the detection of 128 photons on average. Losing just one photon can destroy many edges and substantially increase the generation time. This problem is alleviated by recognizing that the 12 dangling photons of the resource state can be produced deterministically, provided that the highly-connected central ring of 12 qubits is first prepared across 12 sources. Moreover, we exploit local complementation (LC) [^1] LC is a graph operation that corresponds to complementing the neighborhood of a vertex in a graph, and when two graphs can be obtained from one another by LC, the corresponding graph states can be obtained from one another by local Clifford operations [52]. to further reduce the number of edges that need to be established [53]. Using the Graph State Compass (GSC) python package [54], we find that this 12-qubit graph has a minimum-edge representation with just 14 edges, which can be transformed into the desired 12-qubit central ring with four Hadamard gates, see Figure 3c. The minimum-edge representation is hereafter called the LC-graph. This approach works only because the dangling photons can be produced afterwards, since the full 24-photon resource state is already in a minimum-edge representation.
@@
The LC-graph has further useful properties. Most qubits have a low degree of connectivity and so photon loss is less destructive. In addition, it can be constructed using a divide-and-conquer strategy [51, 55] by connecting together identical subgraphs in three stages. Thus, if a photon is lost when generating one subgraph, it will not destroy the others.
@@
Once the LC-graph is complete, it is transformed into the central 12-qubit ring of the resource state using LC. Each source then emits a photon, mapping the prepared graph onto the emitted photons (See Appendix B for details of this process). Since the sources remain entangled to the photonic graph, they can emit a second photon to produce the 12 dangling photons of the final resource state. Finally each spin is measured to decouple it from the graph and reset for the next cycle.
@@ III.1 Figures of Merit
To compare architectures, we define two figures of merit quantifying the performance of an RSG: the resource efficiency $\eta_{R}$ and the transmission efficiency $\eta_{T}$. These metrics are designed to describe RSG architectures that operate on a periodic clock cycle with rate $r$, where a single resource state is produced by the RSG at the end of every cycle with an overall RSG success probability denoted as $p_{\mathrm{rsg}}$. This requirement is important to ensure synchronization between different RSGs and synchronization with the clock cycle of the fusion network required to perform FBQC.
@@ III.1.1 Resource efficiency
To benchmark RSG overhead, a platform-agnostic metric is needed that captures trade-offs across RSG architectures. This metric could encompass factors like the energetic cost of operation, the fabrication cost, or the spatial footprint of the device. Designing an inclusive and comprehensive definition of resource efficiency, as was initiated in [56], is a challenge that should be addressed in a future work. Here, we consider a simple definition that enables the comparison of RSG architectures exploiting spatial and temporal multiplexing schemes.
To quantify the resource efficiency in the spatial domain, we use the ratio of the number of photons in the final resource state $N$ to the number of sources $N_{0}$ used to produce at least one such resource state with a desired probability $p_{\mathrm{rsg}}$. In the temporal domain, we consider the ratio of the RSG clock rate $r$ to the rate $r_{0}$ of the laser pump used to trigger the sources (the internal clock of the RSG). Combining these, we define a spatio-temporal resource efficiency of:
@@ III.1.2 Upper bound on $\eta_{R}$
A spatial MUX does not decrease RSG rate ($r=r_{0}$) and so $\eta_{R}=N/N_{0}$ depends only on the number of sources $N_{0}$. If the architecture must consume an average of $N_{\text{avg}}$ photons to produce one resource state, then a necessary condition to arbitrarily improve the RSG success probability $p_{\mathrm{rsg}}$ toward $1$ is that the $N_{0}$ sources produce a minimum of $N_{\text{avg}}$ photons on average. Thus we have $p_{s}N_{0}\geq N_{\text{avg}}$ which implies $\eta_{R}\leq p_{s}N/N_{\text{avg}}=\tilde{\eta}_{R}$. Note that $p_{s}=1$ for all but the HSPS. This upper bound can be generalized to $r\neq r_{0}$ by noting that an optimal temporal MUX allows one to reduce $N_{0}$ at the cost of reducing $r$ by the same proportion. Hence:
@@
Knowing the upper bound, it is convenient to define the discrepancy between $\eta_{R}$ and $\tilde{\eta}_{R}$, which is directly attributed to the ability for a MUX to use the available photons. This leads to a natural definition of MUX efficiency $\eta_{\text{mux}}=N_{\text{avg}}/(p_{s}N_{0})\cdot(r/r_{0})$ so that
@@
This final expression is convenient as it clearly separates the three contributions to the resource efficiency: the success probability $p_{s}$ that depends only on the source, the MUX efficiency $\eta_{\text{mux}}$ that depends on the multiplexing network, and the efficiency of photon consumption $N/N_{\text{avg}}$ that depends only on the scheme. Note that $\eta_{\mathrm{mux}}$ is dependent on $p_{s}$ and $N_{\mathrm{avg}}$, since the optimal multiplexing scheme, and thus $\eta_{\mathrm{mux}}$, may change depending on the photon generation probability and overall scheme.
The above derivation of $\tilde{\eta}_{R}$ assumes that sources operate continuously for the entire RSG cycle. This is not the case for the RUS module because sources may idle after generating an edge. For example, in Figure 3c, the last operation before LC will be a RUS gate to establish the edge connecting just two of the 12 sources. When this RUS gate is being applied, the 10 other sources must wait without emitting photons, which increases the number of internal cycles per source. Thus $N_{\text{avg}}$ will underestimate the cost in time and overestimate the upper bound on resource efficiency. In this case, the upper bound can instead be computed knowing the average number of internal clock cycles $\tau_{\text{avg}}$ used to produce a resource state:
@@
In the special case where each source attempts to produce a single photon at each internal cycle of the RSG, $N_{\mathrm{avg}}=p_{s}N_{0}\tau_{\mathrm{avg}}$ and we recover Eq. (2).